OpenStudy (anonymous):

Hey guys please help? Medal award :D A sample of radioactive material has been decaying for 5 years. Three years ago (at 2 years ) there were 6.0grams of material left. Now (at 5 years) there are 5.2 grams of material left a. Write two points from this problem b. using the point-ratio y=y1b^x-x1 plug in the two points and solve for the ratio b. c.write an exponential equation in point-ratio form for this situation

OpenStudy (anonymous):

for the two points i believe it is (2,6.0) and (5,5.2) for the point ratio i thnk it's 5.2=6.0b^5-2 c i'm not sure

OpenStudy (anonymous):

Can someone please help..:(

OpenStudy (anonymous):

i believe ur first two are correct but C i have no clue

OpenStudy (anonymous):

@crabby_patty I'm not sure what they mean by exponential point ratio maybe y=ab^x?

OpenStudy (nincompoop):



OpenStudy (anonymous):

since we have the two points what exactly are they looking for?

OpenStudy (anonymous):

(2, 6.0) and (5,5.2)

OpenStudy (nincompoop):

find the common ratio, look at the ratios of consective y-values

OpenStudy (nincompoop):

5.2/6

OpenStudy (anonymous):

and 5/2? or just 5.2/6 i'm just not aware of this concept

OpenStudy (nincompoop):

if you have more than one ratio then you obtain the mean

OpenStudy (mathmale):

The point of this problem is to come up with an EXPONENTIAL model for the relationship between x and y. If we succeed in doing that, then either of the given points will "satisfy" the equation for this model.

OpenStudy (nincompoop):

your ratio determines the exponential decay

OpenStudy (mathmale):

I had some trouble with "y=y1b^x-x1 at first. It took me a while to realize that there should be parentheses around " x - x1 ":\[y=y _{1}b ^{(x-x _{1})}\]

OpenStudy (anonymous):

so would b and c be the same type of equation?

OpenStudy (mathmale):

In other words, the quantity inside parentheses has to be calculated before it's used as the exponent of base b.

OpenStudy (nincompoop):

sorry I misspoke the ratio would be multiplied to the first term 30* (ratio)^x

OpenStudy (mathmale):

@sana1234 " That's right. One and the same equation applies to both Parts a and b.

OpenStudy (anonymous):

@mathmale it's 5.2=6.0b^3?

OpenStudy (nincompoop):

you start with 6

OpenStudy (mathmale):

I'd prefer to see how you obtained your results. Have you checked both of the given points in your equation 5.2=6.0b^3 ?

OpenStudy (anonymous):

@nincompoop 6=5.2b^3?

OpenStudy (nincompoop):

why cube?

OpenStudy (nincompoop):

y = 6(ratio)^x then you look at the proceeding points

OpenStudy (anonymous):

@nincompoop for c? I was just going over my equation for b

OpenStudy (nincompoop):

x-x1 would be applied to the second point you have remember that you are not starting from zero year here. you started with 2nd year so it can mean that on the first year, the gram must be greater than the second year

OpenStudy (mathmale):

@sana1234 I could help you more if you'd share your work. If you truly want to understand how to obtain an exponential model for this situation, you'll need to calculate a and b by substituting the given points {(3,6), (5.5.2)} into the model I gave you,\[y=y _{1}b ^{(x-x _{1})}\]

OpenStudy (nincompoop):

your y1 = 6 not 5.2

OpenStudy (mathmale):

Have you actually done that?

OpenStudy (anonymous):

@mathmale i'm just really confused haha just for clarification for b the equation is 5.2=6.0b^3 and why would it be (3,6) ?

OpenStudy (mathmale):

Because you have two unknown quantities, a and b, and it is your job to find their values. To do that, you take the model given, and then for each point you substitute the x- and y-coordinates into the model. You'll end up with two equations and that is enough info to enable you to calculate a and b.

OpenStudy (mathmale):

going back to what you've just said: the equation is 5.2=6.0b^3 and why would it be (3,6) ? Again, you have two separate points, (3,6) and (5, 5.2). You have ONE model, which I've typed in before. In each case, the point will give you x1 and y1. Substitute those values into the model. Our goal is to find a and b.

OpenStudy (nincompoop):

@mathmale it's (2, 6)

OpenStudy (nincompoop):

decaying for five years. then 3 years ago it means on 2nd year of decay

OpenStudy (mathmale):

What's (2,6)? (2,6) is a point. Was the objective of this problem to find a point?

OpenStudy (anonymous):

@mathmale it's point ratio form

OpenStudy (mathmale):

I understand we're supposed to use that particular form. While I have the background to figure it out, I haven't actually used that form myself. Can either of you find an example in your textbook (if you have one) or online? I learn by example. If you have an example, could you share an image of it with me? There are other exponential models that would apply in this situation; were I able to find an expo model for the 2 given data points, then I could most probably convert that to the point-ratio form.

OpenStudy (anonymous):

just a sec.. I believe i do

OpenStudy (anonymous):

ok i have a really good example

OpenStudy (anonymous):

An exponential decay function in point-ratio form y=y1*b^x-x1 has the horizontal asymptote y=o. because these data approach a long-run value of 1.25, the exponential function must be translated up 1.25 units. To do so replace y with y-1.25. THe coefficient y1 is also a y value, so you must also replace y1 with y1-1.25 in order to account for the translation

OpenStudy (anonymous):

The point-ratio equation is now y-125=(y1-1.25)*b^x-x1. TO find the value of b, select one point for x1 y1 and a second point (not too close) for (x,y). If you choose (10,1.97) and (50,1.36) you will find 1.36-1.25=(1.97-1.25)*b^50-10 -substitute (10,97( for x1 y1 and 50, 1.36 for x,y

OpenStudy (anonymous):

0.11-0.72b^40 subtract b^40=0.11/0.72 "divide both sides by 0.72 b=(0.11/0.72)^1/40 is approximately 0.9541 (raise both sides to the power of 1/40

OpenStudy (mathmale):

Wow, @sana1234, this is really good stuff, some of which is new to me. I have a suggestion: Could you possibly work on something else for a while and let me look over your course material so that I could come to a better understanding of it? Don't want to waste your time while I learn. But, on the other hand, you seem to be making great progress on your own, for which I'm thankful. What do YOU want to do now? Split, and re-convene later, or you go on your own path?

OpenStudy (anonymous):

thanks so much :) i'll be on here i'm actually looking through some other examples haha i have roughly an hour and a half before i have to head to school though again thanks so much!

OpenStudy (mathmale):

OK! I'll be working on this.

OpenStudy (anonymous):

btw for y-1.25=(y1-1.25) *b^x-x1 I forgot to put a point for 1.25

OpenStudy (anonymous):

@mathmale thanks again so much I really appreciate it a lot!

OpenStudy (mathmale):

I'm still learning my way through this, but believe I have an approach that results in my obtaining the common ratio, b. I've obtained the same value for b in several different ways, which increases my optimism that my b is correct. I start with the point-ratio formula\[y=y _{1}*b ^{(x-x _{1})}\] and rewrite it, first as \[y=y _{1}b ^{x}b ^{-x _{1}}\] and then as\[y=\frac{ y _{1} }{ b ^{x _{1}} }b^x\]

OpenStudy (anonymous):

Makes sense :D

OpenStudy (mathmale):

Our goal is to find b. Note that the coefficient of b^x, above, is just a constant. Therefore, this equation could be modeled by \[y=ab^x\] We have 2 points on the curve: (3,6) and (5,5.2). Substituting them into the above equation, \[6=ab^3 \] and\[5.2=ab^5\] Solving these 2 equations simultaneously for b results in \[b^2=\frac{ 5.2 }{ 6 }=\frac{ 52 }{ 60 }=\frac{ 13 }{ 15 }\] So, if my assumptions and math are correct, \[b=\pm \sqrt{\frac{ 13 }{ 15}}\] Hope this helps you to finish your posted problem. Note, again, that there are multiple ways in which to solve this problem. If it happens that you and arrive at the same value for b using different methods, that assures us that we're on right (if different) tracks. Good luck!!

OpenStudy (anonymous):

I believe it' correct :)) it makes sense.. so basically

OpenStudy (anonymous):

@mathmale thank you so much for your time… I honestly really really do appreciate it :)

OpenStudy (mathmale):

My great pleasure! From what I've seen of your work, you know a lot about this kind of math yourself.